Calculate the wavelength of the radiation emitted, when an excited atom of energy band gap 12.43 keV comes back to ground state.


1
Expert's answer
2020-08-31T07:03:51-0400

Solution.

"\Delta E = \frac{h \times c}{\lambda}"

"\lambda = \frac{h \times c}{\Delta E}"

"\lambda = \frac{6.63*10^{-34} \times 3*10^8}{1.99*10^{-15}} = 0.01 \ nm"

Answer:

"\lambda = 0.01 \ nm"